OpenStudy (anonymous):

HELP PLEASE. Which expressions, when evaluated with z = 8, are negative in value? Choose ALL answers that are correct. A. –2z – 14 B. –25 + 4z C. –2z + 15 D. 30 – 3z

OpenStudy (igreen):

Let's do the first option first. -2z - 14 Plug in z = 8: -2(8) - 14 Multiply: -16 - 14 Subtract: -30 So Option 1 is correct. Option 2: -25 + 4z Plug in z = 8: -25 + 4(8) Multiply: -25 + 32 Add: 7 So Option 2 is incorrect. Option 3: -2z + 15 Plug in z = 8: -2(8) + 15 Multiply: -16 + 15 Add: -1 So Option 3 is also correct. Option 4: 30 - 3z Plug in z = 8: 30 - 3(8) Multiply: 30 - 24 Subtract: 6 So Option D is incorrect.

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @OpiGeode

OpenStudy (anonymous):

I am not new to OpenStudy, and I already knew how to do that, thanks..

OpenStudy (igreen):

Lol, it's hard to tell because OS is messed up.

OpenStudy (anonymous):

also A is correct because -2*8=-16-14=-30

OpenStudy (anonymous):

A and C are correct I took the assessment

OpenStudy (anonymous):

Thanks but I asked this question 9+ days ago. I already took the test.